# On computational complexity and average-case hardness of shallow-depth boson sampling

### v1

@@ III Worst-case hardness of output probability estimation
In this section, we find the worst-case #P-hardness of output probability estimation of shallow-depth linear optical circuits in $\mathcal{B}\mathcal{B}^{*}$, for a fixed input and output $\bm{s}$ within a certain additive imprecision. Our worst-case hardness result for the shallow circuit architecture can be represented as follows.
@@
For $M\geq 2N$, approximating the output probability $p_{\bm{s}}(C)$ to within additive error $2^{-O(N)}$ for any $C$ over linear optical circuit architecture $\mathcal{B}\mathcal{B}^{*}$ is #$\rm{P}$-hard in the worst case.
We briefly sketch the proof of our worst-case hardness result for the shallow circuit architecture $\mathcal{B}\mathcal{B}^{*}$. The proof is based on the result by [25], which showed the simulation hardness of exact boson sampling with constant depth linear optical circuits. Specifically, there exist constant-depth linear optical circuits that can simulate an arbitrary given quantum circuit using post-selection. Also, those constant depth circuits can be embedded in our circuit architecture $\mathcal{B}\mathcal{B}^{*}$. Hence, additively approximating the output probability of any quantum circuit can be reduced to additively approximating the output probability of any circuit in $\mathcal{B}\mathcal{B}^{*}$, with imprecision blowup up to the inverse of post-selection probability. Using the fact that the additive approximation of any quantum circuit is #P-hard for certain additive imprecision [34], we can obtain the worst-case hardness of our shallow circuit architecture $\mathcal{B}\mathcal{B}^{*}$.
@@ IV Average-case hardness of output probability estimation
In this section, we prove the average-case #P-hardness of approximating output probabilities of shallow-depth boson sampling within a certain additive imprecision. We focus on $q$-Kaleidoscope circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ with $q=O(1)$, where the gate number $m$ for such architecture is $m=qM\log M$. From the result of Theorem 3, $(\mathcal{B}\mathcal{B}^{*})^{q}$ has the worst-case hardness of output probability approximation, for $q\geq 1$.
@@
Definition 4 (Random circuit ensemble).
Definition 5 (Random collision-free outcome ensemble).
We define $\,\mathcal{G}_{M,N}$ as the uniform distribution over $\binom{M}{N}$ collision-free outcomes of boson sampling with $M$ modes and $N$ photons. Each outcome $\bm{s}\sim\mathcal{G}_{M,N}$ is an $M$-dimensional output configuration vector for the collision-free outcome, such that $\bm{s}=(s_{1},\dots,s_{M})$ where each $s_{i}\in\{0,1\}$ with $\sum_{i=1}^{M}s_{i}=N$.
Using the definitions above, we first state our result on the average-case hardness over the outcome and circuit instances, for shallow-depth linear optical circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ with $q=O(1)$.
@@
The following problem is #$\rm{P}$-hard under a $\rm{BPP}^{\rm{NP}}$ reduction: for any constant $\delta,\eta\geq 0$ with $\delta+\eta<\frac{1}{4}$, on input a random circuit $U\sim\mathcal{H}_{\mathcal{A}}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ and a random outcome $\bm{s}\sim\mathcal{G}_{M,N}$, compute the output probability $p_{\bm{s}}(U)$ within additive imprecision $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$, with probability at least $1-\delta$ over the choice of $U$ for at least $1-\eta$ over the choice of $\bm{s}$.
@@
In the following, we sketch the proof of Theorem 4, by briefly describing the worst-to-average-case reduction process; we leave in Appendix C a detailed proof of Theorem 4. Since our average-case hardness regards both outcomes and circuits, we first describe how to effectively fix the outcome, so that the remaining problem is to establish worst-to-average-case reduction for fixed output probability over random circuit instances. To do so, our strategy is to randomly permute both a given worst-case circuit and a given fixed outcome. That is, we sample random $M$-mode permutation $\bm{P}$ and permute the worst-case circuit $C_{0}$ and the fixed outcome $\bm{s}_{0}$ equally with $\bm{P}$, where the permuted outcome $\bm{s}=\bm{P}\bm{s}_{0}$ now follows the random outcome ensemble $\mathcal{G}_{M,N}$. Then the fixed worst-case output probability $p_{\bm{s}_{0}}(C_{0})$ is equal to $p_{\bm{s}}(\bm{P}C_{0})$, and thus we can obtain the value $p_{\bm{s}_{0}}(C_{0})$ by inferring $p_{\bm{s}}(\bm{P}C_{0})$ via worst-to-average-case reduction over random circuit instances. Here, $p_{\bm{s}}(\bm{P}C_{0})$ now becomes a new worst-case output probability, such that the revised worst-case circuit $C$ is the randomly permuted circuit $\bm{P}C_{0}$, and the revised fixed outcome $\bm{s}$ is the randomly chosen outcome from $\mathcal{G}_{M,N}$.
To establish the worst-to-average-case reduction from the revised worst-case circuit $C$ to the average-case circuit over $\mathcal{H}_{\mathcal{A}}$ for a fixed outcome $\bm{s}$, our strategy is to perturb the circuit from $\mathcal{H}_{\mathcal{A}}$ with the given worst-case circuit $C$ parameterized by a constant $\theta\in[0,1]$. That is, $\theta=0$ corresponds to the average-case distribution $\mathcal{H}_{\mathcal{A}}$ and $\theta=1$ corresponds to the worst-case circuit $C$. Specifically, we choose a perturbation method such that for small enough $\theta$, the success probability of average-case approximation over perturbed random circuits would not have largely deviated from the ideal case ($\theta=0$), and as $\theta$ grows, the perturbed circuit converges to the worst-case circuit $C$. Using such perturbation method and as long as $\theta$ values are small enough, one can obtain the average-case approximate output probability values with high probability over perturbed circuits, parameterized by different values of $\theta$. Also, we can expect that those average-case values contain some information about the worst-case value $p_{\bm{s}}(C)$, depending on the perturbation method and the values of $\theta$. Assuming that worst-case value $p_{\bm{s}}(C)$ can be inferred using the average-case values with small values of $\theta$, one can finally infer the worst-case value $p_{\bm{s}}(C)$, within a certain imprecision determined by the average-case approximation imprecision and the method for the inference.
Therefore, it is crucial to choose a proper perturbation method to establish the worst-to-average-case reduction successfully. Throughout this work, we use the Cayley path for the perturbation, which was employed in Refs. [33, 21, 34] for the hardness proposals of the random circuit sampling.
Definition 6 (Cayley transform [33]).
@@
Using the Cayley transform defined above, we now define the perturbed random circuit distribution.
Definition 7 (Perturbed random circuit ensemble).
Let $\mathcal{A}$ be the circuit architecture with $m$ number of gates. For the given circuit $C_{0}$ in $\mathcal{A}$ with gates $\{G_{i}\}_{i=1}^{m}$, the circuit $U(\theta)$ is defined with each gate of $C_{0}$ replaced by $G_{i}\rightarrow H_{i}(\theta)G_{i}$, where each $H_{i}(\theta)$ is a Cayley transform of independently distributed local Haar random gate $H_{i}$ $(i\in[m])$ parameterized by $\theta\in[0,1]$. We define $\mathcal{H}_{\mathcal{A},\theta}^{C_{0}}$ as the distribution for such $U(\theta)$. Here, the distribution of the $U(0)$ is $\mathcal{H}_{\mathcal{A}}$, and $U(1)=C_{0}$.
Before proceeding, we should make sure that the success probability of average-case approximation over circuits is still large enough after the perturbation, to establish the reduction process successfully. This is evident in the case that the total variation distance over circuits induced by the perturbation is small enough, as the success probability over circuits perturbs by, at most, the total variation distance.
In fact, Ref. [33] proved that total variation distance between $\mathcal{H}_{\mathcal{A}}$ and $\mathcal{H}_{\mathcal{A},\theta}^{C_{0}}$ is small for comparably small perturbation $\theta$.
Lemma 2 (Movassagh [33]).
Let $\mathcal{A}$ be the circuit architecture with $m$ number of gates. For $\theta\ll 1$ and for any circuit $C_{0}$ in $\mathcal{A}$, total variation distance between $\mathcal{H}_{\mathcal{A},\theta}^{C_{0}}$ and $\mathcal{H}_{\mathcal{A}}$ is $O(m\theta)$.
Therefore, by using small $\theta=O(m^{-1})$, one can upper-bound the total variation distance by an arbitrarily small constant, which implies that the success probability of average-case approximation over circuits also perturbs by at most a small constant.
For the worst-to-average-case reduction, we first sample a random circuit $U(\theta)\sim\mathcal{H}_{\mathcal{A},\theta}^{C_{0}}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ and worst-case circuit $C_{0}$. Using $U(\theta)$ with the same random seed $\{H_{i}\}_{i=1}^{m}$ but with different values of $\theta$ satisfying $\theta=O(m^{-1})$, we obtain the average-case approximation of the output probability $p_{\bm{s}}(U(\theta))$ for each $\theta$, which may enable us to infer the worst-case output probability value $p_{\bm{s}}(U(1))=p_{\bm{s}}(C)$.
To investigate the feasibility of the inference of the worst-case value, we examine the behavior of the function $p_{\bm{s}}(U(\theta))$ characterized by the parameter $\theta$. Using Definition 6, we find that $p_{\bm{s}}(U(\theta))$ can be represented as a low-degree rational function in $\theta$.
Lemma 3.
Let $\mathcal{A}$ be the $q$-Kaleidoscope circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ with $m=qM\log M$ number of gates, and $U(\theta)\sim\mathcal{H}_{\mathcal{A},\theta}^{C_{0}}$ for any $C_{0}$ in $\mathcal{A}$. Then for any outcome $\bm{s}$, the output probability $p_{\bm{s}}(U(\theta))$ can be represented as a degree $(4mN,4mN)$ rational funtion in $\theta$.
Proof.
For given circuit unitary matrix $U(\theta)\sim\mathcal{H}_{\mathcal{A},\theta}^{C_{0}}$ with $C_{0}$ composed of $\{G_{i}\}_{i=1}^{m}$ gates, one can decompose $U(\theta)$ with $m=qM\log M$ product of unitary matrices, such that each matrix element of $U(\theta)$ can be represented as
$\left[U(\theta)\right]_{j,k}=\sum_{l_{1}=1}^{M}\sum_{l_{2}=1}^{M}\cdots\sum_{l_{m-1}=1}^{M}U_{j,l_{1}}^{(1)}U_{l_{1},l_{2}}^{(2)}\cdots U_{l_{m-1},k}^{(m)},$ (6)
where each $U^{(i)}$ denotes an $M$-dimensional unitary matrix, with a single gate unitary matrix $H_{i}(\theta)G_{i}$ applied to the modes participating in the gate and identity for the rest of the modes. For example, if the $i$th gate $H_{i}(\theta)G_{i}$ is a two-mode gate between the first two modes, $U^{(i)}$ is a block diagonal matrix of $H_{i}(\theta)G_{i}$ and identity matrix, namely, $U^{(i)}=H_{i}(\theta)G_{i}\bigoplus I_{M-2}$.
For circuit architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ which is composed of only two-mode gates, matrix elements of $U^{(i)}$ can be represented as degree $(2,2)$ rational functions in $\theta$, where the common denominator for the elements is given by $q_{i}(\theta)$, defined in Eq. (4) but with index $i$ appended for $i$th random gate $H_{i}(\theta)$. Using reduction to the common denominator for all of the $m$ gates, $\left[U(\theta)\right]_{j,k}$ can be represented as $(2m,2m)$ rational function in $\theta$ with the common denominator $\prod_{i=1}^{m}q_{i}(\theta)$; note that it does not change with the indices $j,k$.
From Eq. (1), the output probability $p_{\bm{s}}(U(\theta))$ has the form of
$\begin{split}p_{\bm{s}}(U(\theta))&=\left|\sum_{\sigma\in S_{N}}\prod_{j=1}^{N}\left[U(\theta)_{\bm{s},\bm{t}}\right]_{\sigma_{j},j}\right|^{2},\end{split}$ (7)
where $\bm{t}$ is an input configuration vector, and $S_{N}$ is $N$-mode permutation group. One can easily check that the common denominator for $\prod_{j=1}^{N}\left[U(\theta)_{\bm{s},\bm{t}}\right]_{\sigma_{j},j}$ is $[\prod_{i=1}^{m}q_{i}(\theta)]^{N}$, and it does not change with permutation $\sigma$. Let us define $Q(\theta)=[\prod_{i=1}^{m}|q_{i}(\theta)|^{2}]^{N}$, which is a degree $4mN$ polynomial in $\theta$. Then $Q(\theta)$ serves as the common denominator for the output probability. Hence, the output probability can be represented as $p_{\bm{s}}(U(\theta))=\frac{P(\theta)}{Q(\theta)}$, with $P(\theta)$ also a degree $4mN$ polynomial in $\theta$. ∎
We are now ready to turn to the proof of Theorem 4, i.e., the average-case hardness of the shallow-depth boson sampling. We prove that for high probability over $\bm{s}\sim\mathcal{G}_{M,N}$, well-approximating output probability $p_{\bm{s}}(U)$ with high probability over $U\sim\mathcal{H}_{\mathcal{A}}$ for the shallow-depth architecture $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q}$ is #P-hard under a $\rm{BPP}^{\rm{NP}}$ reduction.
@@
See Appendix C

### v2

@@ III Worst-case hardness of output probability estimation
In this section, we find the worst-case #P-hardness of output probability estimation of shallow-depth linear optical circuits in the circuit architecture $\mathcal{B}\mathcal{B}^{*}$ defined in Definition 3, for a fixed input $\bm{t}$ and a fixed collision-free output $\bm{s}_{0}$, to within a certain additive imprecision. Our worst-case hardness result for our shallow linear optical circuit architecture can be represented as follows.
@@
Given $M=\Omega(N)$, for a fixed collision-free outcome $\bm{s}_{0}$, approximating an output probability $p_{\bm{s}_{0}}(C_{0})$ to within additive imprecision $2^{-O(N)}$ for any $C_{0}$ over linear optical circuit architecture $\mathcal{B}\mathcal{B}^{*}$ is #$\rm{P}$-hard in the worst case.
We briefly sketch the proof of our worst-case hardness result for the shallow circuit architecture $\mathcal{B}\mathcal{B}^{*}$. The proof is based on the result by [33], which showed the simulation hardness of exact Boson Sampling with constant-depth linear optical circuits. Specifically, as shown in [33], there exists a constant-depth linear optical circuit that can simulate an arbitrary given quantum circuit using post-selection. To extend this result to our case, we show that the constant-depth worst-case circuit in [33] can be embedded in our circuit architecture $\mathcal{B}\mathcal{B}^{*}$. Combining these arguments, additively approximating the output probability of any quantum circuit can be reduced to additively approximating the output probability of any circuit in $\mathcal{B}\mathcal{B}^{*}$, with imprecision blowup up to the inverse of post-selection probability. Using the fact that the additive approximation of any quantum circuit is #P-hard for certain additive imprecision [42], we can obtain the worst-case hardness of our shallow circuit architecture $\mathcal{B}\mathcal{B}^{*}$.
@@ IV Average-case hardness of output probability estimation
In this section, we introduce our average-case hardness result of approximating output probabilities of shallow-depth Boson Sampling. We focus on $q$-Kaleidoscope circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q}$ in Definition 3 with $q=O(1)$, whose circuit depth is $D=2q\log M$ and gate number is $m=qM\log M$. By Theorem 3, $\mathcal{B}\mathcal{B}^{*}$ has the worst-case hardness, and thus $(\mathcal{B}\mathcal{B}^{*})^{q}$ with $q\geq 1$ also has the worst-case hardness by adding trivial gates (identity gates) over the worst-case circuit in $\mathcal{B}\mathcal{B}^{*}$.
@@
We now formally state our average-case hardness result for shallow-depth linear optical circuit architecture $(\mathcal{B}\mathcal{B}^{*})^{q\geq 4}$ with $q=O(1)$.
@@
The following problem is #$\rm{P}$-hard under a $\rm{BPP}^{\rm{NP}}$ reduction: for any constant $\delta,\eta\geq 0$ with $\delta+\eta<\frac{1}{4}$, on input an $O(\log N)$-depth random circuit $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$ and a random collision-free outcome $\bm{s}\sim\mathcal{G}_{M,N}$, compute the output probability $p_{\bm{s}}(U)$ within additive imprecision $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$, with probability at least $1-\delta$ over the choice of $U$ for at least $1-\eta$ over the choice of $\bm{s}$.
We sketch the proof of Theorem 4 in the following by briefly describing the worst-to-average-case reduction process; we leave in Appendix E a detailed proof of Theorem 4. Now, we clarify how to establish the reduction: The goal is to well-estimate worst-case output probability $p_{\bm{s}_{0}}(C_{0})$ for a fixed collision-free outcome $\bm{s}_{0}$ and worst-case circuit $C_{0}$ in $(\mathcal{B}\mathcal{B}^{*})^{q\geq 1}$ (Theorem 3), given oracle access to well-estimated values $p_{\bm{s}}(U)$ for average-case collision-free outcomes over $\bm{s}\sim\mathcal{G}_{M,N}$ and average-case circuits over $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$ (in other words, $U$ is a randomly chosen circuit over $(\mathcal{B}\mathcal{B}^{*})^{q\geq 4}$).
@@
First, since our average-case argument regards both outcomes and circuits, we begin by describing how to fix the outcome in establishing the worst-to-average-case reduction. The remaining problem is then to establish a “circuit-level” worst-to-average-case reduction, from a worst-case circuit to randomly chosen average-case circuits for a fixed output probability, similarly to previous hardness results of random circuit sampling [40, 41, 28, 42, 43]. To fix the outcome, our strategy is to randomly permute both a given worst-case circuit $C_{0}$ in $(\mathcal{B}\mathcal{B}^{*})^{q\geq 1}$ and a given collision-free outcome $s_{0}$. That is, we sample random $M$-mode permutation $\bm{P}_{0}$ uniformly, and permute the worst-case circuit $C_{0}$ and the fixed outcome $\bm{s}_{0}$ in Theorem 3 equally with $\bm{P}_{0}$. Then, the permuted outcome $\bm{s}=\bm{P}_{0}\bm{s}_{0}$ essentially follows the random outcome ensemble $\bm{s}\sim\mathcal{G}_{M,N}$, and the permuted circuit $\bm{P}_{0}C_{0}$ is in $(\mathcal{B}\mathcal{B}^{*})^{q\geq 2}$ by Lemma 1. Note that the fixed worst-case output probability $p_{\bm{s}_{0}}(C_{0})$ is now equivalent to $p_{\bm{s}}(\bm{P}_{0}C_{0})$, and thus our new goal is to estimate $p_{\bm{s}}(\bm{P}_{0}C_{0})$ value, given estimated average-case values of $p_{\bm{s}}(U)$.
Next, we describe how to estimate $p_{\bm{s}}(\bm{P}_{0}C_{0})$ given average-case estimation values $p_{\bm{s}}(U)$ for $U\sim\mathcal{H}_{\mathcal{A}}\mathcal{P}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$. For the permutation part $\mathcal{P}$ in the average-case circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$, we newly sample an $M$-mode random permutation $\bm{P}_{1}\sim\mathcal{P}$ again, and using the sampled $\bm{P}_{1}$ we set the circuit $\bm{P}_{0}C_{0}\bm{P}_{1}^{-1}$ as a “revised” worst-case circuit in $(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$. For the local random circuit part $\mathcal{H}_{\mathcal{A}}$ in the average-case circuit ensemble $\mathcal{H}_{\mathcal{A}}\mathcal{P}$, we sample a random circuit $V$ from $\mathcal{H}_{\mathcal{A}}$ and perturb each gate of $V$ by each gate of $\bm{P}_{0}C_{0}\bm{P}_{1}^{-1}$ parameterized by a constant $\theta\in[0,1]$ using Cayley transform method, and obtain a perturbed circuit $V(\theta)$ (see Appendix D for details). Here, by Cayley transform, $V(\theta)$ lies on a continuous path over unitary matrices interpolating $V(0)$ and $V(1)$, where $\theta=0$ corresponds to the random circuit drawn from local random circuit ensemble $V(0)\sim\mathcal{H}_{(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}}$ and $\theta=1$ corresponds to the worst-case circuit $V(1)=\bm{P}_{0}C_{0}\bm{P}_{1}^{-1}$. Notably, $V(0)\bm{P}_{1}$ follows the average-case circuit distribution $\mathcal{H}_{\mathcal{A}}\mathcal{P}$ with $\mathcal{A}=(\mathcal{B}\mathcal{B}^{*})^{q\geq 3}$, and thus, one can expect that the distribution for $V(\theta)\bm{P}_{1}$ for sufficiently small $\theta$ would not be largely deviated from the average-case circuit distribution. Accordingly, one can obtain well-estimated values of $p_{\bm{s}}(V(\theta)\bm{P}_{1})$ for sufficiently small $\theta$ with high probability, given oracle access to the average-case output probabilities (i.e., Theorem 4).
Importantly, we find that $p_{\bm{s}}(V(\theta)\bm{P}_{1})$ can be expressed as a finite-degree rational function $\frac{P(\theta)}{Q(\theta)}$ for $\theta\in[0,1]$, where $Q(\theta)$ is efficiently computable. Hence, for sufficiently small $\theta$ values, one can also obtain the well-estimated values for the finite degree polynomial $P(\theta)$, by multiplying the well-estimated values of $p_{\bm{s}}(V(\theta)\bm{P}_{1})$ with $Q(\theta)$. Then, given the estimated values of $P(\theta)$ for different $\theta$ values, one can use polynomial interpolation for the polynomial $P(\theta)$, and obtain the estimate of $P(1)=p_{\bm{s}}(V(1)\bm{P}_{1})Q(1)$. Therefore, one can finally infer the value $p_{\bm{s}}(V(1)\bm{P}_{1})$ by multiplying $Q(1)^{-1}$ to the estimate of $P(1)$, which corresponds to the worst-case output probability $p_{\bm{s}}(V(1)\bm{P}_{1})=p_{\bm{s}}(\bm{P}_{0}C_{0}\bm{P}_{1}^{-1}\bm{P}_{1})=p_{\bm{s}}(\bm{P}_{0}C_{0})=p_{\bm{s}_{0}}(C_{0})$, concluding the worst-to-average-case reduction.
To sum up, for the worst-case output probability $p_{\bm{s}_{0}}(C_{0})$ in Theorem 3 and the average-case output probability $p_{\bm{s}}(U)$ in Theorem 4, the overall reduction process can be schematized by
$p_{\bm{s}_{0}}(C_{0})=p_{\bm{s}}(\bm{P}_{0}C_{0})=p_{\bm{s}}(\bm{P}_{0}C_{0}\bm{P}_{1}^{-1}\bm{P}_{1})\quad\xleftarrow[\begin{subarray}{c}V\rightarrow\bm{P}_{0}C_{0}\bm{P}_{1}^{-1}\\ \text{Cayley transform}\\ \text{and}\\ \text{Polynomial interpolation}\end{subarray}]{}\quad p_{\bm{s}}(V\bm{P}_{1})=p_{\bm{s}}(U)\approx\mathcal{O}(\bm{s},U),$ (2)
@@
[…]
$C_{0}:\text{Worst-case circuit in }(\mathcal{B}\mathcal{B}^{*})^{q\geq 1}$
[…]
So far, we have outlined the proof of Theorem 4, describing how to establish worst-to-average-case reduction. In Appendix D, we introduce in detail the Cayley transform, the circuit perturbation method for the worst-to-average-case reduction in our work. Subsequently, in Appendix E, we provide a detailed proof of Theorem 4, including the imprecision analysis we have missed out in the main text (i.e., why the average-case imprecision level $\epsilon=2^{-O(N^{\gamma+1}(\log N)^{2})}$ is necessary to achieve the worst-case imprecision level $2^{-O(N)}$).
@@
See Appendix E
It is worth remarking that, although we have concentrated on the dilute regime such that $\gamma\geq 2$, our average-case hardness result does not rule out the case $21\leq\gamma<$ 2 because it is derived from the worst-to-average-case reduction and our worst-case hardness result allows $M=\Omega(N)$. However, since our average-case hardness result only considers collision-free outcomes, it does not directly lead to the sampling hardness result (i.e., quantum advantage claim for shallow-depth Boson Sampling) by only increasing the allowed additive imprecision for the average-case hardness because collision-free outcomes are expected to occupy a small portion of outcomes when $21\leq\gamma<$ 2. Therefore, to lead to the sampling hardness argument when $21\leq\gamma<$ 2, our average-case hardness result should be extended to also contain collision outcomes for the average-case outcome instances (i.e., instead of $\mathcal{G}_{M,N}$, now uniformly choose over all possible outcomes). For a more complete analysis, in Appendix G, we present a promising approach that can be extended to the average-case hardness result that includes collision outcomes.